a shipping box measures 14 1/2 inches long by 9 1/8 inches wide by 8 11/16 inches high. What is the volume of the box in cubic feet
95.79 ft3
0.67 ft3
0.02 ft3
0.78 ft3

Answer by checkley71(8403)

14.5*9.125*8.6875=1,149.4648 in^3 or seeing as there are 12*12*12=1,728 In 6 in one qubic foot we have 1,149.4648/1,728=.6651995 cubic feet or rounding we get .67 ft^3